A spring oscillator is designed with a mass of 0.1480.148 kg. It operates while immersed in a damping fluid, selected so that the oscillation amplitude will decrease to 1.00% of its initial value in 5.845.84 s. Find the required damping constant for the system.
Answer #1

A = A0 e(-lt)

0.01=e(-l*5.84)

ln100=5.84*l

l=0.7885

we know that l=b/2m

so b=2ml

damping constant = 2ml = 2*0.148*0.7885 = 0.2153 kg/sec Answer

feel free to drop comment if you have any doubt.